OpenStudy (anonymous):

How can an expression written in either radical form or rational exponent form be rewritten to fit the other form? HELP ME PLEASE I WILL GIVE MEDALS AND A FAN

OpenStudy (goformit100):

Sir/Ma'am Please show your working

OpenStudy (anonymous):

@goformit100 it's just a response question there are no numbers or anything i just need to explain how

OpenStudy (anonymous):

I forgot this a bit but if i'm not mistaken the number inside the radical is the numerator and the number outside the radical sign is the denominator in the rational exponent form, if thats right then you just do the same thing with the exponent to find the radical form. \[3\sqrt{x^2}=x \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

Example^

OpenStudy (anonymous):

Thank you!!! @JoshR1213